### v1

# Diffuse Neutrino Background from Magnetorotational Stellar Core Collapses

Abstract
A statistically significant detection of the diffuse supernova neutrino background (DSNB) is around the corner. To this purpose, we assess the contribution to the DSNB of magnetorotational collapses of massive stars, relying on a suite of state-of-the-art three-dimensional neutrino-magnetohydrodynamic simulations. We find that neutrinos from magnetorotational core collapses boost the high-energy tail of the DSNB spectrum, similar to what is expected from neutrino-driven black hole-forming collapses. The latest data from the Super-Kamiokande Collaboration can already exclude that more than $13\%$ of all collapsing massive stars undergo magnetorotational collapses under optimistic assumptions. A DSNB detection at $3\sigma$ could take place up to $4$ yr earlier at Super-Kamiokande-Gadolinium or JUNO if the fraction of magnetorotational collapses should be larger than $10\%$. Fascinatingly, if the fraction of magnetorotational stellar collapses should be larger than $7\%$, Hyper-Kamiokande could measure such a fraction at $3\sigma$ after $20$ yr of DSNB data taking. The combination of DSNB and electromagnetic data has the potential to resolve the degenerate contributions from magnetorotational and neutrino-driven black hole forming collapses, providing crucial insight on the properties of the population of collapsing massive stars.
I Introduction
On average, every second, a core-collapse supernova (CCSN) explodes somewhere in the Universe releasing $\sim 10^{53}$ erg in the form of neutrinos and antineutrinos of all flavors. The cumulative flux of neutrinos emitted from all CCSNe in the Universe over cosmological distances is the Diffuse Supernova Neutrino Background (DSNB) [1, 2, 3, 4, 5, 6, 7]. The DSNB is expected to be the dominant contribution to the overall diffuse emission of neutrinos from known sources between $10$ and $25$ MeV [8].
To date, the theoretical modeling of the DSNB is hampered by our incomplete understanding of the CCSN rate [9] and the initial mass function [10], the rate of collapses forming a black hole (BH)  [11, 12], flavor conversion in the supernova (SN) core [13], the metallicity evolution of galaxies [14, 15], binary stellar interactions [16], as well as the neutron star properties [17]; we refer the interested reader to Ref. [17] for an overview of the relevant uncertainties. Moreover, Type Ia SNe [18], population III supermassive stars [19, 20, 21], and accretion flows [22, 23] can also contribute to the diffuse neutrino emission in the same energy region.
The DSNB is the key observational target of the Super-Kamiokande-Gadolinium project [24, 25]. While the DSNB signal has not been detected with high confidence yet, the first results from the Super-Kamiokande-Gadolinium project are extremely encouraging [24, 25, 26, 27].
So far, most DSNB searches have focused on the detection of electron antineutrinos via inverse beta decay (i.e., $\bar{\nu}_{e}+p\to e^{+}+n$). Whereas this approach remains the most promising one [28, 24, 26, 27], a measurement of the DSNB in all flavors, and for neutrino and antineutrinos, is desirable. In this regard, the liquid Argon neutrino telescope DUNE should improve existing limits on $\nu_{e}$ [29, 30]; detectors based on the neutrino-nucleus coherent elastic scattering, such as DARWIN and RES-NOVA, could probe the non-electron flavors [31, 32, 33].
In this paper, we investigate the contribution to the diffuse neutrino emission from magnetorotational core collapses of massive stars, leading to protomagnetars and spinars.  [^1] Note that we distinguish between magnetorotational core collapses leading to the formation of a protomagnetar or a spinar (if the central compact object is a BH). The term “spinar” was first introduced in Ref. [34] in the context of active galactic nuclei, but it is now commonly adopted to refer to a fast rotating magnetized relativistic object, in quasi-equilibrium due to the balance between the centrifugal and gravitational forces, which undergoes a two-stage collapse, first to a protoneutron star and, subsequently, to form a BH [35, 36, 37, 38]. Magnetorotational core collapses are considered to be the progenitors of superluminous SNe and certain gamma-ray bursts [39, 40]. To this purpose, we model the neutrino emission from the (transiently stable) protomagnetars relying on three-dimensional neutrino-magnetohydrodynamic simulations of the same models studied in 2D in Ref. [40]. Since the rate of protomagnetars and spinars is unknown, we keep it as a free parameter and investigate whether the neutrino emission from magnetorotational core collapses could lead to a DSNB contribution which is degenerate with respect to the uncertainties in the CCSN rate and fraction of neutrino-driven BH-forming collapses.
Our work is organized as follows. In Sec. II, we introduce the suite of models adopted to compute the DSNB. Section III presents our fiducial DSNB model from CCSNe and neutrino-driven BH-forming collapses as well as explore the contribution from protomagnetars and spinars. Section IV presents a comparison between our theoretical prediction and the model-independent limits on the $\bar{\nu}_{e}$ flux from Super-Kamiokande [27]. A large fraction of magnetorotational core collapses would enhance the near-future prospects for detecting the DSNB at Super-Kamiokande-Gadolinium and JUNO, as shown in Sec. V. There, we also explore the sensitivity of Hyper-Kamiokande loaded with Gadolinium to measure the fraction of protomagnetars and spinars contributing to the DSNB. Section VI highlights the relevance of combining the DSNB data with electromagnetic observations to overcome the astrophysical uncertainties currently plaguing the DSNB. Finally, we conclude in Sec. VII.
II Neutrino driven and magnetorotational core collapses: Neutrino emission properties
In this section, we introduce the neutrino emission properties for successful neutrino-driven CCSNe and neutrino-driven BH-forming models, as well as the protomagnetar and spinar models adopted to reproduce the population of collapsing massive stars.
II.1 Core-collapse supernova and black hole forming collapse models
[…]
Figure 1 displays the temporal evolution of the neutrino emission properties (for $\nu_{e}$, $\bar{\nu}_{e}$, and $\nu_{x}$, where the latter denotes $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{\tau}$, and $\bar{\nu}_{\tau}$) for the neutrino-driven collapse models introduced above. One can see that the duration of the neutrino signal is shorter for the neutrino-driven BH-forming model, with neutrino luminosity slightly larger than the one of the neutrino-driven CCSN models.
Figure 1: Temporal evolution of the neutrino emission properties for one representative protomagnetar model with mass of $17~{}M_{\odot}$ (1D equivalent neutrino emission properties of model A17, black lines), the spinar model with mass 26 $M_{\odot}$ (1D equivalent neutrino emission properties of model of model A26, gray lines), two 1D neutrino-driven CCSNe models with mass of $11.2~{}M_{\odot}$ and $27~{}M_{\odot}$ (red and orange lines, respectively), and a 1D neutrino-driven BH-forming collapse with $40~{}M_{\odot}$ (model s4027b2 characterized by a short accretion phase before prompt BH formation, green lines). The upper and lower panels display the luminosity and mean energy, for $\nu_{e}$, $\bar{\nu}_{e}$, and $\nu_{x}$, respectively. The protomagnetar and spinar models exhibit larger average energies for the non-electron flavors due to neutrino decoupling occurring at higher densities than for the neutrino-driven CCSN models.
II.2 Protomagnetar and spinar models
In order to model the neutrino emission from protomagnetars we rely on 3D simulations combining special relativistic magnetohydrodynamics and a spectral two-moment (M1) neutrino transport [43, 40]. The simulations use the SFHo hadronic equation of state [44] and include the most relevant neutrino-matter reactions (absorption, emission, and scattering by nucleons and nuclei, scattering off electrons and positrons, electron-positron pair annihilation and nucleonic bremsstrahlung). The models are the 3D equivalent of the 2D ones presented in Ref. [40]. We call protomagnetar a post-collapsed core that develops magnetar magnetic field strengths $\gtrsim\mathcal{O}(10^{15})$ G and has an associated magnetar phenomenology as described in, e.g., Refs. [45, 46]; if the collapse into a BH occurs a few seconds after core bounce, we refer to our models as spinars. Note that we do not consider the contribution to the neutrino background from the resulting BH accretion disk, which was investigated in Ref. [22].
We rely on seven models with mass at the zero age main sequence time: $M_{\mathrm{ZAMS}}=5~{}M_{\odot}$, $8~{}M_{\odot}$, $13~{}M_{\odot}$, $17~{}M_{\odot}$, $20~{}M_{\odot}$, $26~{}M_{\odot}$, and $30~{}M_{\odot}$, obtained from the progenitors presented in Ref. [39]. These models consider enhanced rotational mixing, leading to chemically homogeneous evolution. While a detailed description of the dynamics and explosion properties of these 3D models will be presented elsewhere, here we provide a brief summary.
All cores produce delayed explosions at times between $\sim 0.3$ and $\sim 1.3$ s after bounce. Shock revival is powered by a combination of neutrino heating and magnetorotational stresses, with a higher contribution of the latter leading to explosions with energies above $E_{\mathrm{exp}}\gtrsim 10^{51}$ erg. Unlike the 2D models presented in Ref. [40], only the 3D models with the larger masses ($M_{\mathrm{ZAMS}}\gtrsim 26M_{\odot}$; models A26 and A30 in the nomenclature of that reference) have clear signs of BH formation, leading to spinars. The formation of the BH does not happen promptly after core bounce, but, at least, after $1.5$ s. This is the case of model A26, for example. For more massive models, more than $2.5$ s are needed for the protomagnetars to collapse. A common feature of these high-mass models is that the protoneutron star mass grows (due to non-spherical accretion) above the limit set by the equation of state relatively soon after bounce. High-mass progenitors thus form protomagnetars for a relatively short period of time, and then they may turn into BH-forming events or even collapsars at later times, as anticipated in Refs. [47, 48]. Lower mass progenitors (including models A05, A08, A13, A17, and A20) do not display signatures of BH formation; the protoneutron star mass remains below the maximum mass set by the equation of state throughout the computed evolution (for some models, such as A17, the evolution lasts for more than $3.5$ s after bounce). These lower mass progenitor models are more likely to form standard protomagnetars that survive the SN explosion. It is noteworthy that the progenitor stars modeled in Ref. [39] were predicted to produce either superluminous SNe (if $M_{\textrm{ZAMS}}\lesssim 13M_{\odot}$), protomagnetar-driven SNe ($M_{\textrm{ZAMS}}=17M_{\odot}$, $20M_{\odot}$), or collapsars ($M_{\textrm{ZAMS}}\gtrsim 26M_{\odot}$). However, the limited simulated time for our 3D models does not allow to distinguish whether the protomagnetar models lead to superluminous SNe or protomagnetar-driven explosions, so far. Nevertheless, the distribution of specific angular momentum in spinars (see, e.g., Fig. 2 of Ref. [40]) suggests promising prospects for the formation of an accretion disc, and hence, the potential for a collapsar scenario.
The M1 neutrino transport evolves the spectral energy and momentum densities, $J(\vec{r},E)$ and $\vec{H}(\vec{r},E)$, of $\nu_{e}$, $\bar{\nu}_{e}$, and $\nu_{x}$ as functions of position ($\vec{r}\,$) and particle energy ($E$). As the momentum density is equal to the energy flux of the neutrinos, we can extract the luminosity directly from its radial component by integrating it at an extraction radius, fixed at $r_{\mathrm{ext}}=500$ km, over the solid angle: $L_{\nu}(E)=\int_{\Omega}\mathrm{d}\Omega\,r_{\mathrm{ext}}^{2}J^{r}(r_{\mathrm{ext}},E)$. The number density of the neutrinos and its flux can be obtained by dividing the spectral energy and momentum density by the particle energy, $\{N(\vec{x},E),\vec{F}_{N}(\vec{x},E)\}={1}/{E}\{J(\vec{x},E),\vec{H}(\vec{x},E)\}$. We then compute the neutrino mean energies via an integration over the energy spectrum: $\langle\varepsilon_{\nu}\rangle=\langle\int\mathrm{d}E\,J(\vec{x},E)\rangle_{\Omega,r_{\mathrm{ext}}}/\langle\int\mathrm{d}E\,N(\vec{x},E)\rangle_{\Omega,r_{\mathrm{ext}}}$, where $\langle...\rangle_{\Omega,r_{\mathrm{ext}}}$ indicates an average over the solid angle at the extraction radius.
Figure 1 shows the angle-averaged (1D equivalent) neutrino emission properties for the representative protomagnetar model with mass $M=17M_{\odot}$ (model A17; note that the sudden termination of the neutrino signal at about $4$ s corresponds to the end of the simulation time for this model and not to BH formation) and the spinar model with mass 26 $M_{\odot}$ (model A26; the signal terminates about $1.5$ s after core bounce due to BH formation). These models have neutrino luminosity $\text{L}_{\nu}\gtrsim 10^{52}$ erg s-1 for all flavors, with the typical neutrino emission lasting for several seconds. The neutrino mean energies for the different flavors exhibit a characteristic hierarchy: $\langle\varepsilon_{\nu_{e}}\rangle<\langle\varepsilon_{\bar{\nu}_{e}}\rangle<\langle\varepsilon_{\nu_{x}}\rangle$. The large mean energy of the non-electron flavor neutrinos (which exceeds $20$ MeV for a significant part of the emission) is due to the compact protoneutron star, that facilitates the neutrino decoupling at densities higher than in the neutrino-driven CCSN models. The small offset (of a few MeV) in the heavy-neutrino mean energies is produced by the development of steep density and temperature gradients at the typical densities where the transition region from diffusion to free streaming is located. As a result, the flux factors (i.e., the ratio of flux-to-neutrino energy) grow over a narrow radial range. The region where $\nu_{x}$’s decouple from matter and the other neutrino flavors coincides with a sharp drop of the gas temperature, such that even a small radial difference has an impact on the mean energies of the non-electron neutrino flavors. The detailed treatment of the pair processes may have also some (small) impact on the computed heavy-neutrino mean energy.
III Diffuse neutrino background
In this section, after outlining the approach employed to model the DSNB, we compute the contribution of neutrino-driven CCSN explosions, neutrino-driven BH-forming collapses, protomagnetars, and spinars to the diffuse neutrino emission.
III.1 Contribution to the diffuse neutrino background from neutrino-driven core-collapse supernovae and black hole-forming collapses
The diffuse neutrino background from neutrino-driven CCSNe and neutrino-driven BH forming collapses is given by:
$\Phi_{\text{CCSN}}(E)=$ $c\int_{8M_{\odot}}^{125M_{\odot}}\textrm{d}M$
×$\,\times\int_{z=0}^{z_{\text{max}}}\text{d}z\frac{\mathcal{R}(z,M)}{H(z)}F_{\nu}(M,E(1+z))\,,$ (1)
where $c$ denotes the speed of light, $z$ is the redshift with $z_{\text{max}}=5$ (see below), $\mathcal{R}(z,M)$ is the CCSN rate, $F_{\nu}(M,E(1+z))$ is the time-integrated neutrino energy spectrum of a CCSN of mass $M$, and the factor $1/H(z)$ accounts for the expansion history of the Universe.
Neutrino energetics.
The time-integrated neutrino energy spectrum, $F_{\nu}(M,E)$, is computed as
[…]
where we integrate over the duration of neutrino emission the product of the luminosity ($\mathrm{L}_{\nu}(E,t)$), the spectral energy distribution ($\phi^{0}_{\nu}(E,t)$ parametrized following Refs. [49, 50]), and the neutrino mean energy ($\langle\varepsilon_{\nu}(t)\rangle$).
The top panel of Fig. 2 shows the $\bar{\nu}_{e}$ DSNB spectrum computed following Eq. 1 and assuming that all collapsing massive stars behave as the $11.2~{}M_{\odot}$, or $27~{}M_{\odot}$ CCSN models, or the $40~{}M_{\odot}$ neutrino-driven BH-forming model (i.e., we compute the DSNB assuming that each of these models would be representative of the whole core collapse population in order to highlight the differences in the neutrino emission). One can see that, because of the energetic and long-lasting emission from protomagnetars (cf. $17~{}M_{\odot}$ model), their flux is the largest with the high-energy tail as soft as that of BH-forming CCSNe. Similarly, the luminous neutrino emission from spinars results in the second largest flux (after protomagnetars, cf. $26M_{\odot}$ model) with a soft high-energy tail comparable to the one of protomagnetars.
Figure 2: Top panel: Electron-antineutrino DSNB flux (Eq. 1) as a function of the observed neutrino energy, assuming that the $11.2~{}M_{\odot}$ or the $27~{}M_{\odot}$, the $40~{}M_{\odot}$ model (dashed, dotted, and dashed-dotted gray lines), the protomagnetar model with mass of $17~{}M_{\odot}$ (dashed orange line), or the spinar model with mass of $26~{}M_{\odot}$ (dashed green line) are representative of the population of collapsing massive stars, respectively. Flavor conversion at the source is neglected. Bottom panel: Same as top panel, but now taking into account different CCSN and magnetorotational models according to the initial mass function (cf. Fig. 3) and for different values of the fraction of magnetorotational collapses ($f_{\text{MR}}$, colored lines). The dark gray band shows the fiducial DSNB model and the light gray one takes into account complete flavor conversion at the source.
Core-collapse supernova rate.
The bulk of the DSNB mainly comes from neutrinos emitted from core collapses at $z\lesssim z_{\text{max}}$ [51]. For these redshifts, the expansion rate of the Universe is well described by $H(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}$, with the local expansion rate being $H_{0}=70$ km s-1 Mpc-1 [52], the dark-energy density $\Omega_{\Lambda}=0.7$, and the matter density $\Omega_{m}=0.3$ [53].
We parametrize the CCSN rate in terms of the initial mass function (IMF) and the star formation history $\dot{\rho}_{*}(z)$:
$\mathcal{R}(z,M)=\frac{\text{IMF}(M)}{\int_{0.1M_{\odot}}^{125M_{\odot}}\text{d}MM\times\text{IMF}(M)}\dot{\rho}_{*}(z)\ ,$ (3)
for collapsing stars with mass $\gtrsim 8M_{\odot}$. For what concerns the star formation history, we adopt the following piecewise parametrization [54]:
$\dot{\rho}_{*}(z)=\dot{\rho}_{0}\Bigg{[}$ $(1+z)^{\alpha\eta}$
$+\left(\frac{1+z}{B}\right)^{\beta\eta}+\left(\frac{1+z}{C}\right)^{\gamma\eta}\Bigg{]}^{1/\eta}\,,$ (4)
with $\eta=-10$. The constants $B$ and $C$ are given by
$B=(1+z_{1})^{1-\alpha/\beta}\,,$ (5)
$C=(1+z_{1})^{(\beta-\alpha)/\gamma}(1+z_{2})^{1-\beta/\gamma}\,.$ (6)
Given our poor knowledge of the CCSN rate, which constitutes the largest uncertainty plaguing the DSNB [13, 54], we consider a range of values for $\dot{\rho}_{0}$, $\alpha$, $\beta$, and $\gamma$ as summarized in Table 1, and we set $z_{1}=1$ and $z_{2}=4$.
Table 1: Numerical values of the parameters in the upper, fiducial, and lower analytical fits of the star formation rate from Ref. [54], assuming the Salpeter initial mass function [55].
Ranges $\dot{\rho}_{0}$ $\alpha$ $\beta$ $\gamma$
Upper $0.0213$ $3.6$ $-0.1$ $-2.5$
Fiducial $0.0178$ $3.4$ $-0.3$ $-3.5$
Lower $0.0142$ $3.2$ $-0.5$ $-4.5$
The analytical fits of $\dot{\rho}_{*}(z)$ depend on the assumed initial mass function [54]. We consider for the latter the Salpeter initial mass function [55]: $\text{IMF}\propto(M/M_{\odot})^{-2.35}$. Employing non-universal initial mass functions would affect the CCSN rate (i.e., $\dot{\rho}_{0}$) with a negligible impact on the DSNB [56, 10].
Fraction of neutrino-driven black hole-forming collapses.
Another unknown in the DSNB modeling is the fraction of non-rotating (or mildly-rotating) massive stars collapsing into BHs. To take this into account, we allow for the fraction of neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}$) to vary, following Ref. [29], and restrict the range to $f_{\nu\text{BH}}<0.47$. This choice is motivated by the findings of Ref. [17], which considers a suite of parameterized 1D neutrino-driven engines for single-star and helium-star progenitors (i.e., $0.4170.178<f_{\nu\text{BH}}<$ 0.417).
Neutrino flavor conversion.
The DSNB is expected to be affected by neutrino flavor conversions [13]. However, due to the large uncertainties in the modeling of neutrino quantum kinetics in the core of SNe [57, 58, 6], we rely on a conservative approach and consider two extreme cases: (i) no flavor conversion, or (ii) full flavor conversion (i.e., $\nu_{e}$ or $\bar{\nu}_{e}$ emerge from the source as $\nu_{x}$ or $\bar{\nu}_{x}$, swapping completely their flavor). We stress that none of these assumptions is realistic, yet they should serve as guidance to assess the impact of flavor conversion uncertainties. Conversely, flavor conversion of neutrinos en route to Earth is not subject to large uncertainties [59]. To model this, we consider the best-fit values of the neutrino mixing parameters for normal mass ordering from Ref. [60].
The dark gray band in the bottom panel of Fig. 2 represents the $\bar{\nu}_{e}$ component of our “fiducial” DSNB flux model, obtained for $0.470<f_{\nu\text{BH}}<$ 0.47, $0.02130.0142<\dot{\rho}_{0}<$ 0.0213, $3.63.2<\alpha<$ 3.6, $-0.5<\beta<-0.1$, and $-4.5<\gamma<-2.5$, as well as no flavor conversion. The light gray band of Fig. 2 instead corresponds to the case of full flavor conversion. We assume the values for the fiducial parametrization of the CCSN rate (cf. Table 1) and $f_{\nu\text{BH}}=0.3$, which is an average value according to Ref. [17]. In order to make this plot, we consider each of the core collapse models from our suite as representative of the population for a certain mass range (cf. the top panel of Fig. 3). The slope of the tail of the spectrum becomes less steep as the fraction of neutrino-driven BH-forming collapses increases, since the emission from this source class is more energetic.
III.2 Contribution to the diffuse neutrino background from protomagnetars and spinars
Assuming that our protomagnetar model A17 or our spinar model A26 are representative of the population of massive stellar collapses, the top panel of Fig. 2 shows the corresponding $\bar{\nu}_{e}$ fluxes.  [^2] We have verified that the neutrino spectral energy distributions of our suite of magnetorotational collapse models follow the parametrization of Refs. [49, 50]. One can see that the $\bar{\nu}_{e}$ spectra from protomagnetars and spinars exhibit an energy tail with slope comparable to the one of the neutrino-driven BH-forming collapse model, but peaking at $E\simeq 10$ MeV in the source frame.
When taking into account the contribution of protomagnetars and spinars, the total diffuse neutrino emission is
$\Phi_{\text{diffuse}}=(1-f_{\text{MR}})\Phi_{\nu\text{CCSN}}+f_{\text{MR}}\Phi_{\text{MR}}\ ,$ (7)
where $\Phi_{\nu\text{CCSN}}$ is the diffuse flux from neutrino-driven CCSNe and neutrino-driven BH forming collapses (Eq. 1), while $\Phi_{\text{MR}}$ represents the contribution from protomagnetars and spinars, and $f_{\text{MR}}$ is the fraction of magnetorotational collapses assumed to be constant over redshift.
For the initial mass function of magnetorotational stellar core collapses, we adopt the Salpeter one, but extending from $5M_{\odot}$ to $125M_{\odot}$. The lower minimum mass that we consider for magnetorotational collapses ($5M_{\odot}$) is motivated by the fact that the chemically homogeneous evolution of the models under consideration results in larger carbon-oxygen cores compared to those in non-rotating, standard core-collapse progenitors. These cores provide conditions conducive to the development of subsequent nuclear burning stages, even below the non-rotating threshold of approximately $8M_{\odot}$. Consequently, chemically homogeneous models in the mass range $[5M_{\odot},8M_{\odot}]$ may also undergo core collapse instead of evolving into carbon-oxygen white dwarfs. The lower panel in Fig. 3 illustrates how the magnetorotational models introduced in Sec. II are employed to model the magnetorotational core collapse population, where for each band in the initial mass function we have considered the correspondent protomagnetar or spinar model.
We have also considered two extreme initial mass functions, $\text{IMF}_{\rm MR,1}\propto(M/M_{\odot})^{-2.7}$ and $\text{IMF}_{\text{MR, 2}}\propto(M/M_{\odot})^{-1}$, following Refs. [61] and [62], respectively. We find that the DSNB is negligibly affected by the choice of the IMF, in agreement with the findings of Ref. [10]; we, thus, only report our findings for the Salpeter mass function.
Figure 3: Top panel: Initial mass function for neutrino-driven CCSNe and BH forming collapses as a function of the ZAMS progenitor mass. Bottom panel: Same as the top panel, but for the magnetorotational core collapses. The neutrino-driven core collapse and the protomagnetar/spinar populations extend up to $125M_{\odot}$, although not explicitly shown in the plots. The vertical dashed lines highlight the ZAMS mass of the models in our suite. The color bands correspond to the range of masses for which we consider each model of our suite as representative. In parenthesis, we indicate if the remnant compact object is a neutron star (NS) or a BH, following Refs. [6, 41] and the 3D models for the magnetorotational collapses adopted in this work (note that this setup fixes the fraction of spinars to $f_{\text{BH-MR}}\simeq 0.11$).
The diffuse neutrino emission from protomagnetars and spinars is obtained analogously to Eq. 1. It is worth noticing that the prefactor $f_{\text{MR}}$ in Eq. 7 accounts for the different normalization in the initial mass function of CCSNe and the one of magnetars resulting from the different lower mass limit ($8\,M_{\odot}$ or $5\,M_{\odot}$) adopted for each of the two cases. In addition, in the computation of the contribution of neutrinos from magnetorotational collapses to the DSNB, the uncertainties in the star formation history and flavor conversion are taken into account as described in Sec. III.1.
The lower panel of Fig. 2 shows the contribution of different fractions of magnetorotational collapses to the DSNB, one can see that the tail of the latter increases with the fraction of protomagnetars and spinars. It is worth noticing that the magnetorotational collapse models predict an emission of non-electron flavor neutrinos and antineutrinos with larger mean energies than the electron-flavor ones. Hence, uncertainties due to flavor conversion in the source can potentially affect the spectral shape of the related diffuse neutrino emission even more than in CCSNe.
[…]
Figure 4: Electron-antineutrino DSNB for $f_{\text{MR}}=0.03$ (green) and $f_{\text{MR}}=0.1$ (orange), including the contribution from neutrino-driven CCSNe and BH forming collapses and in the absence of flavor conversion. The gray band corresponds to the fiducial DSNB prediction (i.e., for $f_{\text{MR}}=0.1$, uncertainties on the star formation history as from Table 1, and $f_{\nu\text{BH}}<0.47$). The solid gray line corresponds to the fiducial cosmic rate (cf. Table 1) and $f_{\nu\text{BH}}=0.3$. The error bars represent the model-independent upper limits from Super-Kamiokande on a flux of $\bar{\nu}_{e}$’s of astrophysical origin [27]. The neutrino emission from magnetorotational collapses enhances the high-energy tail of the DSNB spectrum, boosting its detection prospects.
IV Constraining the fraction of magnetorotational collapses with the Super-Kamiokande-Gd data
Several neutrino telescopes carry out DSNB searches, including KamLAND [64], Borexino [65], SNO [66] and Super-Kamiokande [67, 25]. Recently, the Super-Kamiokande Collaboration reported their most stringent limits on a model-independent $\bar{\nu}_{e}$ flux of astrophysical origin using data from the phases VI and VII, with different levels of Gadolinium doping [27]. We compare such limits to the DSNB flux expected at Super-Kamiokande-Gadolinium computed as
$\bar{\Phi}(E_{i},E_{f})=\frac{\int_{E_{i}}^{E_{f}}\sigma_{\text{IBD}}(E)\Phi_{\text{diffuse}}(E)\text{d}E}{\int_{E_{i}}^{E_{f}}\sigma_{\text{IBD}}(E)\text{d}E}\ ,$ (8)
where $\sigma_{\text{IBD}}$ is the cross-section for inverse beta decay [68, 69]. For the sake of simplicity, we assume an energy-independent detection efficiency and perfect energy resolution [70].
The large number of reactor antineutrinos below $10$ MeV and the atmospheric background above $\sim 30$ MeV constrain the neutrino detection window for the DSNB [4, 5]. The model-independent results reported by the Super-Kamiokande Collaboration account for these backgrounds and assume an energy-independent flux per energy bin. As a consequence, when comparing with our theoretical predictions, we unfold the energy-dependence of the detection cross section–see Eq. (8).
Figure 5 compares the limits reported by the Super-Kamiokande Collaboration with our DSNB forecast. We consider the fiducial DSNB model ($f_{\text{MR}}=0$, gray band), the DSNB computed for $f_{\text{MR}}=0.03$ (top panel, dark green bands) and $f_{\text{MR}}=0.1$ (bottom panel, dark orange bands) and also scenarios with full flavor conversion at the source (see Sec. III, light green and light orange bands, respectively). In all cases, our forecast for the DSNB signal falls below the current Super-Kamiokande limits. However, if $f_{\text{MR}}\gtrsim 0.13$ and full flavor conversion would take place for the highest local star formation rate, the Super-Kamiokande-Gadolinium upper limits would already be incompatible with the DSNB model. Note that, in order to allow for a direct quantitative comparison of our model to the data, the energy binning used in Fig. 5 is analogous to the one adopted by the Super-Kamiokande-Gadolinium Collaboration [27].
Figure 5: Electron-antineutrino DSNB flux binned in energy for $f_{\text{MR}}=0.03$ (top panel) and $f_{\text{MR}}=0.1$ (bottom panel). The gray bands correspond to the fiducial DSNB model (i.e., obtained considering uncertainties on the star formation history as from Table 1), $f_{\nu\text{BH}}<0.47$, and no flavor conversion at the source. The dark shaded colored bands correspond to the DSNB obtained for $f_{\text{MR}}=0.03$ and $f_{\text{MR}}=0.1$. Light shaded bands are obtained for full flavor conversion at the source. The black dots and lines represent the $\bar{\nu}_{e}$ model-independent limits obtained by the Super-Kamiokande Collaboration [27]. Current Super-Kamiokande data do not allow to constrain $f_{\text{MR}}\lesssim 0.1$.
V Detection prospects and future sensitivity
In this section, we discuss the impact of a non-zero fraction of magnetorotational core collapses in the near-future DSNB detection prospects at Super-Kamiokande-Gadolinium [71] and JUNO [72]. We also forecast the sensitivity of Hyper-Kamiokande-Gadolinium [73, 74] and DUNE [75] to the fraction of magnetorotational collapses. The detector characteristics that we consider for these four neutrino telescopes are summarized in Table 2. This table also includes the number of signal events for $f_{\text{MR}}=0.1$ and $f_{\text{MR}}=0$ as well as the number of background events.
V.1 Near term detection prospects
The increase in Gadolinium doping in Super-Kamiokande is expected to result in a larger event statistics also for what concerns the DSNB [76, 77]. To illustrate the short-term potential of Super-Kamiokande-Gadolinium, we investigate the exposure that one would need to detect the DSNB and infer the fraction of magnetorotational collapses.
[…]
The expected number of signal events for a neutrino observatory is
$S=N_{t}\tau\xi\int\sigma_{\text{IBD}}(E)\Phi_{\text{diffuse}}(E)\text{d}E\,,$ (9)
where $N_{t}$ is the number of targets for a given detector volume, $\tau$ is the exposure time, and $\xi$ is the energy-independent detection efficiency. The integral over the neutrino energy is carried out for the energy window of interest (see Table 2)
For a fiducial volume of $22.5$ kT, $65\%$ detection efficiency, and the backgrounds reported for the Super-Kamiokande-VI and Super-Kamiokande-VII runs [27], we compute the ratio between the signal events ($S$) and the Poissonian statistical fluctuations of the total number of events, i.e. signal plus background ($B$). For our baseline analysis, the fraction of magnetorotational collapses is the only free parameter for a fixed star formation history (i.e., for a given set of values $\dot{\rho}_{0}$, $\alpha$, $\beta$ and $\gamma$; see Table 1)–whereas the CCSN rate ( Eq. 4) and the fraction of neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}=0.3$) are assumed to be known.
In the absence of systematic uncertainties, $S/\sqrt{S+B}=3$ corresponds to reject the background-only hypothesis with a significance of $3\sigma$. Precisely, the top panel of Fig. 6 shows the region where $S/\sqrt{S+B}=3$, in the plane spanned by the exposure and the fraction of magnetorotational collapses. The light-shaded band in the top panel of Fig. 6 corresponds to the case of full flavor conversion at the source. Similarly, JUNO is expected to start taking data in late 2024 [78]. With a fiducial volume of $17$ kT, detection efficiency of $80\%$, and backgrounds as described in Ref. [79], its DSNB detection prospects are comparable to the ones from Super-Kamiokande doped with Gadolinium, as visible from the bottom panel of Fig. 6. For both Super-Kamiokande-Gadolinium and JUNO, a non-zero $f_{\text{MR}}$ would result in a positive signal at neutrino observatories within less than a decade of data taking. In particular, a non-zero fraction of magnetorotational collapses could result in the DSNB detection within the next 66668888 yr or earlier.
We stress that the evolution of the sensitivity of Super-Kamiokande depends on the Gadolinium-enrichment phases and other experimental aspects such as background rejection and neutron tagging [28, 24, 26]. Changes in the experimental setup in JUNO may also modify the results presented above quantitatively.
Table 2: Fiducial volume, detection efficiency ($\xi$), and optimal energy window for JUNO [79], Super-Kamiokande-Gadolinium (Super-K-Gd) [27], DUNE [75], and Hyper-Kamiokande-Gadolinium (Hyper-K-Gd) [73]. We also indicate the energy-integrated number of DSNB events for $f_{\text{MR}}=0$ and $f_{\text{MR}}=0.1$, and the number of background events per $100$ kT and year.
Experiment
Fiducial
volume
[[[[kt]]]]
Detection
efficiency
Energy
window
[[[[MeV]]]]
DSNB events
($f_{\text{MR}}=0$)
[…]
DSNB events
($f_{\text{MR}}=0.1$)
[…]
Background
events
[…]
JUNO $17$ $0.7$ 9.49.49.49.431.431.431.431.4 $11.0$ $14.3$ $3.2$
Super-K-Gd $22.5$ $0.65$ 9.49.49.49.431.431.431.431.4 $16.9$ $21.9$ $26.1$
DUNE $40$ $1$ 19.019.019.019.031.031.031.031.0 $1.3$ $1.8$ $0.6$
Hyper-K-Gd $374$ $0.67$ 9.49.49.49.431.431.431.431.4 $17.5$ $22.6$ $18.0$ ($10.5$  [^3] This value corresponds to the scenario in which neutral current backgrounds induced by atmospheric neutrinos are negligible.)
Importantly, the combination of data from different neutrino observatories–affected by different systematics and backgrounds–could further advance the detection of the DSNB and hint at the existence of a population of magnetorotational collapses contributing to the DSNB. For instance, combining Super-Kamiokande-Gadolinium and JUNO would result in an increase of the statistical significance of an eventual DSNB detection, as evident from the expected number of events summarized in Table 2.
V.2 Long term detection prospects
A better assessment on the contribution from protomagnetars and spinars to the DSNB is expected from the upcoming water Cherenkov neutrino telescope Hyper-Kamiokande [73]. We consider an experimental configuration with two tanks of $187$ kT each and detection efficiency of $67\%$, resulting from adding $\sim 0.1\%$ of GdCl3 to water (cf. Table 2 for a summary on the expected number of events). Among the different backgrounds expected in Hyper-Kamiokande, neutral-current atmospheric background events are particularly relevant [80, 81], and the possibility of reducing them to negligible level is under investigation [82].
Figure 7 shows the sensitivity to the fraction of magnetorotational collapses as a function of the exposure for the DSNB model of Sec. III, with and without neutral current atmospheric backgrounds. This corresponds to the sensitivity to distinguish the DSNB prediction for a non-zero fraction of magnetorotational collapses from a DNSB signal with $f_{\text{MR}}=0$, i.e. we consider a ratio:
$\frac{S(f_{\text{MR}})-S(f_{\text{MR}}=0)}{\sqrt{S(f_{\text{MR}}=0)+B}}=3\,.$ (10)
One can see that a fraction of magnetorotational collapses smaller than $7\%$ could be measured at $3\sigma$ after $20$ yr of data taking, assuming that the fraction of neutrino-driven BH-forming collapses is known and the neutral current atmospheric background is efficiently tagged.
Figure 7: Hyper-Kamiokande sensitivity at $3\sigma$ to the fraction of magnetorotational core collapses ($f_{\text{MR}}$) as a function of the exposure. The solid line corresponds to negligible neutral-current backgrounds induced by atmospheric neutrinos. The dashed line considers the fiducial backgrounds from Ref. [83]. The dark shaded band represents the impact of the uncertainties on the CCSN rate and the fraction of neutrino-driven BH-forming collapses. After $20$ yr exposure, Hyper-Kamiokande-Gadolinium will be sensitive to $f_{\rm{MR}}\gtrsim 7\%$.
The upcoming liquid Argon neutrino experiment DUNE [75], being sensitive to the $\nu_{e}$ component, could also add positively to this picture, although with smaller statistics (cf. Table 2). We estimate that a ratio $S/\sqrt{S+B}=3$ (i.e. to reject the background-only hypothesis at $3\sigma$ level) could only be achieved in $10$ yr at this detector, if $f_{\text{MR}}\gtrsim 0.27$–assuming $100\%$ detection efficiency and the backgrounds from Ref. [83]. The difference between the mean energies of electron and non-electron flavor neutrinos is larger than the one for antineutrinos. This could potentially disentangle some of the existing uncertainties in the picture, especially for what concerns the flavor conversion physics and the fraction of neutrino-driven BH-forming collapses [29].
VI Complementary electromagnetic constraints
Assuming magnetorotational core collapses are distributed following the star formation rate, their fraction and the one of neutrino-driven BH-forming collapses are degenerate in the DSNB signal. In both cases, the larger the population, the higher would be the high-energy tail of the DSNB spectrum. To illustrate this, we assume that the true fractions of neutrino-driven BH-forming collapses and magnetorotational collapses are $f_{\nu\text{BH}}=0.3$ and $f_{\text{MR}}=0$ and compute the expected number of DSNB events after $20$ yr of data taking at Hyper-Kamiokande-Gadolinium, in the absence of flavor conversion. Figure 8 shows the region of the parameter space (i.e. values of $f_{\text{MR}}$ and $f_{\nu\text{BH}}$) compatible with our hypothesis. To guide the eye, we also display the allowed fraction of neutrino-driven BH-forming collapses from the numerical simulations of Ref. [17] as a green-shaded contour. Note that limits on BH formation could also be inferred from searches with the Large Binocular Telescope on disappearing luminous stars [84]; we select from Ref. [84] the constraints obtained assuming only one failed SN: $f_{\nu\text{BH}}=0.16^{+0.23}_{-0.12}$, as not all of the luminous stars in their sample may be neutrino-driven BH-forming collapses. These limits are displayed in Fig. 8 (shaded blue region).
Figure 8: Present and future constraints from neutrino (DSNB) electromagnetic data and numerical simulations on the fraction of magnetorotational collapses ($f_{\text{MR}}$) and neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}$). The red-shaded region corresponds to the projected sensitivity of Hyper-Kamiokande-Gadolinium after $20$ yr of data, taking into account the uncertainties on the CCSN rate. To guide the eye, the black dot marks our fiducial DSNB model. The orange band represents the observed rate of successful explosions constrained by the Palomar facility [85]. The green-shaded region displays the interval of $f_{\nu\text{BH}}$ extracted from simulations of the CCSN population [17]. Additional constraints from the search for disappearing luminous stars from the Large Binocular Telescope [84] are shown in blue. The vertical dashed line corresponds to the upper limits on the observed rate of stripped-envelope SNe (SESNe) at the Palomar transient facility [85]. The neutrino and electromagnetic data and numerical simulations can potentially provide complementary constraints on $f_{\text{MR}}$ and $f_{\nu\text{BH}}$.
While we have focused on neutrinos, combining DSNB and electromagnetic data could help to break the degeneracy in the fraction of magnetorotational collapses plaguing the DSNB data. The Palomar Transient Facility has measured the total rate of observed SNe–both CCSN and stripped-envelope SNe–to be [85]:
$\mathcal{R}^{\text{obs}}_{\text{Palomar, total}}=9.10^{+1.56}_{-1.27}\times 10^{-5}\text{yr}^{-1}\text{Mpc}^{-3}\,,$ (11)
for $H_{0}=70$ $\cdots$-1Mpc-1 and at an average redshift of $\langle z\rangle=0.028$. The fraction of BH-forming collapses as well as the one of protomagnetars and spinars could be inferred relying on the Palomar data, as shown in Fig. 8.
The Palomar Transient Facility has also determined the local rate of stripped-envelope SNe to be [85]:
$\mathcal{R}^{\text{obs}}_{\text{Palomar, SESN}}=2.41^{+0.81}_{-0.64}\times 10^{-5}\text{yr}^{-1}\text{Mpc}^{-3}\,.$ (12)
One can then constrain the fraction of magnetorotational collapses by requiring that the observable rate of protomagnetars and spinars predicted in this work is smaller than the rate of stripped SNe observed by the Palomar Transient Facility. This would provide an upper limit, since it would rely on the assumption that all stripped SNe are linked to our suite of magnetorotational collapses, i.e., $f_{\rm MR}<\mathcal{R}^{\text{obs}}_{\text{Palomar, SESN}}/\mathcal{R}^{\text{obs}}_{\text{Palomar, total}}\sim 0.26$.
VII Conclusions and outlook
As the detection of the DSNB approaches, it is timely to address whether the diffuse neutrino emission from magnetorotational core collapses, leading to the formation of protomagnetars or spinars, could contribute to the DSNB. We rely on a suite of seven 3D special relativistic neutrino-magnetohydrodynamic simulations with different ZAMS mass, inspired by the 2D equivalent models presented in Ref. [40], to investigate the related neutrino properties and the associated contribution to the DSNB. Our models predict large neutrino luminosities $\gtrsim 10^{52}$ ergs-1 for several seconds and large mean energies for non-electron neutrinos and antineutrinos ($\langle\varepsilon_{\nu_{x}}\rangle\gtrsim 20$ MeV). Such neutrino properties suggest that magnetorotational core collapses could provide a non-negligible contribution to the high-energy DSNB tail (above $5$ MeV).
[…]
Given its larger size, the upcoming Hyper-Kamiokande-Gadolinium neutrino telescope could measure a fraction of magnetorotational collapses larger than $7\%$ at $3\sigma$ after $20$ yr of data-taking, if meanwhile we can infer the fraction of (neutrino-driven) BH-forming collapses and pin down the uncertanties on the CCSN rate through electromagnetic observations. This measurement would further benefit from improved background reduction and detection efficiency.
The observational signatures of magnetorotational collapses in the DSNB signal are entangled with the ones linked to neutrino-driven BH-forming collapses. However, complementary electromagnetic probes, e.g., searches for disappearing luminous stars [86, 87, 84], could help to break such degeneracies. Moreover, previous work [39, 40] suggests that a subset of our suite of magnetorotational models could harbor progenitors of long gamma-ray bursts (LGRBs). Assuming that the magnetorotational progenitors with ZAMS mass $M\gtrsim 15M_{\odot}$ lead to GRB jets [39] and that magnetorotational core collapses are the only sources of LGRBs, one could compare the rate of LGRBs with our findings to break the DSNB degeneracies. Similarly, the connection between these magnetorotational collapses and superluminous SNe [39] could also be explored. Such multi-messenger approach would greatly benefit from reduced uncertainties on the star formation history [88] as well as progress in stellar evolution models [89, 90, 91, 92], and our understanding of angular momentum transport in massive stars [93, 94, 95, 96].
With this work, we set a roadmap for an improved comprehension of the sources that can contribute to the DSNB, particularly protomagnetars and spinars. While existing data are still plagued by astrophysical uncertainties, the combination of upcoming neutrino and electromagnetic measurements will be crucial to learn more about the properties of the population of magnetorotational collapses in our Universe.
Acknowledgements.
In Copenhagen, this project has received support from the Villum Foundation (Project No. 13164), the Danmarks Frie Forskningsfond (Project No. 8049-00038B), the Carlsberg Foundation (CF18-0183) and the Deutsche Forschungsgemeinschaft through Sonderforschungbereich SFB 1258 “Neutrinos and Dark Matter in Astro- and Particle Physics” (NDM). Part of this work was performed at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. In Valencia, this work was supported from the grant PID2021-127495NB-I00, funded by MCIN/AEI/10.13039/501100011033 and by the European Union under “NextGenerationEU” as well as “ESF: Investing in your future”, the Astrophysics and High Energy Physics program of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union “NextGenerationEU” (PRTR-C17.I1), and the Prometeo excellence program grant CIPROM/2022/13 funded by the Generalitat Valenciana. MO acknowledges support from the Spanish Ministry of Science via the Ramón y Cajal programme (RYC2018-024938-I). The Tycho supercomputer hosted at the SCIENCE HPC Center at the University of Copenhagen and computing time granted via the Red Española de Supercomputación (grants AECT-2022-3-0016, AECT-2023-1-0009, AECT-2023-3-0019, AECT-2024-1-0013) at the supercomputers MareNostrum-4 and MareNostrum-5 of the Barcelona Supercomputing Centre as well as on the hosts Tirant (grants AECT-2022-1-0003, AECT-2023-1-0001) and Lluisvives of the Servei d’Informàtica of the Universitat de València were used for supporting the numerical simulations presented in this work.

### v2

@@
 A DSNB detection at $3\sigma$ could take place up to $4$ yr earlier at Super-Kamiokande-Gadolinium or JUNO if the fraction of magnetorotational collapses should be larger than $10\%$.
-Fascinatingly, if the fraction of magnetorotational stellar collapses should be larger than $7\%$, Hyper-Kamiokande could measure such a fraction at $3\sigma$ after $20$ yr of DSNB data taking.
-The combination of DSNB and electromagnetic data has the potential to resolve the degenerate contributions from magnetorotational and neutrino-driven black hole forming collapses, providing crucial insight on the properties of the population of collapsing massive stars.
+Fascinatingly, if the fraction of magnetorotational stellar collapses should be larger than $19\%$ ($13\%$), Hyper-Kamiokande could measure such a fraction at $3\sigma$ after ($10$ yr) $20$ yr of DSNB data taking.
+The combination of DSNB and electromagnetic data has the potential to resolve the degenerate contributions from magnetorotational and neutrino-driven black hole-forming collapses, providing crucial insight on the properties of the population of collapsing massive stars.
 I Introduction
@@
 In this regard, the liquid Argon neutrino telescope DUNE should improve existing limits on $\nu_{e}$ [29, 30]; detectors based on the neutrino-nucleus coherent elastic scattering, such as DARWIN and RES-NOVA, could probe the non-electron flavors [31, 32, 33].
-In this paper, we investigate the contribution to the diffuse neutrino emission from magnetorotational core collapses of massive stars, leading to protomagnetars and spinars.
-[^1] Note that we distinguish between magnetorotational core collapses leading to the formation of a protomagnetar or a spinar (if the central compact object is a BH).
-The term “spinar” was first introduced in Ref.
-[34] in the context of active galactic nuclei, but it is now commonly adopted to refer to a fast rotating magnetized relativistic object, in quasi-equilibrium due to the balance between the centrifugal and gravitational forces, which undergoes a two-stage collapse, first to a protoneutron star and, subsequently, to form a BH [35, 36, 37, 38].
+In this paper, we investigate the contribution to the diffuse neutrino emission from magnetorotational core collapses of massive stars, i.e.
+from the collapse of rapidly rotating stars for which the rotation and magnetic fields are relevant.
+We differentiate between two magnetorotational collapse scenarios according to the nature of their remnant: a protomagnetar or a spinar.
+We refer to rapidly rotating protoneutron stars with strong magnetic fields as protomagnetars.
+We adopt the term spinar for a fast rotating magnetized relativistic object, in quasi-equilibrium due to the balance between the centrifugal and gravitational forces, which undergoes a two-stage collapse, first to a protoneutron star and, subsequently, to form a BH [34, 35, 36, 37]  [^1] The term “spinar” was first introduced in Ref.
+[38] in the context of active galactic nuclei.
 Magnetorotational core collapses are considered to be the progenitors of superluminous SNe and certain gamma-ray bursts [39, 40].
@@
 Figure 1 shows the angle-averaged (1D equivalent) neutrino emission properties for the representative protomagnetar model with mass $M=17M_{\odot}$ (model A17; note that the sudden termination of the neutrino signal at about $4$ s corresponds to the end of the simulation time for this model and not to BH formation) and the spinar model with mass 26 $M_{\odot}$ (model A26; the signal terminates about $1.5$ s after core bounce due to BH formation).
-These models have neutrino luminosity $\text{L}_{\nu}\gtrsim 10^{52}$ erg s-1 for all flavors, with the typical neutrino emission lasting for several seconds.
+These models have neutrino luminosity $\text{L}_{\nu}\gtrsim 10^{52}$ $\cdots$-1 for all flavors, with the typical neutrino emission lasting for several seconds.
 The neutrino mean energies for the different flavors exhibit a characteristic hierarchy: $\langle\varepsilon_{\nu_{e}}\rangle<\langle\varepsilon_{\bar{\nu}_{e}}\rangle<\langle\varepsilon_{\nu_{x}}\rangle$.
@@
 [49, 50].
-One can see that the $\bar{\nu}_{e}$ spectra from protomagnetars and spinars exhibit an energy tail with slope comparable to the one of the neutrino-driven BH-forming collapse model, but peaking at $E\simeq 10$ MeV in the source frame.
+One can see that the $\bar{\nu}_{e}$ spectra from protomagnetars and spinars exhibit an energy tail with slope comparable to the one of the neutrino-driven BH-forming collapse model.
+However, for magnetorotational core collapses, the energy distributions peak at $E\simeq 10$ MeV in the source frame.
 When taking into account the contribution of protomagnetars and spinars, the total diffuse neutrino emission is
@@
 V.1 Near term detection prospects
-The increase in Gadolinium doping in Super-Kamiokande is expected to result in a larger event statistics also for what concerns the DSNB [76, 77].
+The increase in Gadolinium doping in Super-Kamiokande has resulted in a larger event statistics also for what concerns the DSNB [76, 77].
 To illustrate the short-term potential of Super-Kamiokande-Gadolinium, we investigate the exposure that one would need to detect the DSNB and infer the fraction of magnetorotational collapses.
@@
 The integral over the neutrino energy is carried out for the energy window of interest (see Table 2)
-For a fiducial volume of $22.5$ kT, $65\%$ detection efficiency, and the backgrounds reported for the Super-Kamiokande-VI and Super-Kamiokande-VII runs [27], we compute the ratio between the signal events ($S$) and the Poissonian statistical fluctuations of the total number of events, i.e.
+For a fiducial volume of $22.5$ kT, $55\%$ detection efficiency  [^3] The detection efficiency is the result of a $75\%$ neutron-capture efficiency for a $0.03\%$ of Gd-loaded water and a $74\%$ selection efficiency [25]., and the backgrounds reported for the Super-Kamiokande-VI and Super-Kamiokande-VII runs [27], we compute the ratio between the signal events ($S$) and the Poissonian statistical fluctuations of the total number of events, i.e.
 signal plus background ($B$).
-For our baseline analysis, the fraction of magnetorotational collapses is the only free parameter for a fixed star formation history (i.e., for a given set of values $\dot{\rho}_{0}$, $\alpha$, $\beta$ and $\gamma$; see Table 1)–whereas the CCSN rate ( Eq.
+For our baseline analysis, the fraction of magnetorotational collapses is the only free parameter for a fixed star formation history (i.e., for a given set of values $\dot{\rho}_{0}$, $\alpha$, $\beta$ and $\gamma$; see Table 1)–whereas the CCSN rate (see Eq.
 4) and the fraction of neutrino-driven BH-forming collapses ($f_{\nu\text{BH}}=0.3$) are assumed to be known.
@@
 […]
-JUNO $17$ $0.7$ 9.49.49.49.431.431.431.431.4 $11.0$ $14.3$ $3.2$
-Super-K-Gd $22.5$ $0.65$ 9.49.49.49.431.431.431.431.4 $16.9$ $21.9$ $26.1$
+JUNO $17$ $0.8$ 9.49.49.49.431.431.431.431.4 $11.0$ $14.3$ $3.2$
+Super-K-Gd $22.5$ $0.55$ 9.49.49.49.431.431.431.431.4 $14.3$ $18.6$ $26.1$
 DUNE $40$ $1$ 19.019.019.019.031.031.031.031.0 $1.3$ $1.8$ $0.6$
-Hyper-K-Gd $374$ $0.67$ 9.49.49.49.431.431.431.431.4 $17.5$ $22.6$ $18.0$ ($10.5$  [^3] This value corresponds to the scenario in which neutral current backgrounds induced by atmospheric neutrinos are negligible.)
+Hyper-K-Gd $187$ $0.55$ 9.49.49.49.431.431.431.431.4 $14.3$ $18.6$ $18.0$ ($10.5$  [^4] This value corresponds to the scenario in which neutral current backgrounds induced by atmospheric neutrinos are negligible.)
 Importantly, the combination of data from different neutrino observatories–affected by different systematics and backgrounds–could further advance the detection of the DSNB and hint at the existence of a population of magnetorotational collapses contributing to the DSNB.
@@
 V.2 Long term detection prospects
-A better assessment on the contribution from protomagnetars and spinars to the DSNB is expected from the upcoming water Cherenkov neutrino telescope Hyper-Kamiokande [73].
-We consider an experimental configuration with two tanks of $187$ kT each and detection efficiency of $67\%$, resulting from adding $\sim 0.1\%$ of GdCl3 to water (cf.
-Table 2 for a summary on the expected number of events).
-Among the different backgrounds expected in Hyper-Kamiokande, neutral-current atmospheric background events are particularly relevant [80, 81], and the possibility of reducing them to negligible level is under investigation [82].
+[…]
 Figure 7 shows the sensitivity to the fraction of magnetorotational collapses as a function of the exposure for the DSNB model of Sec.
@@
 we consider a ratio:
-$\frac{S(f_{\text{MR}})-S(f_{\text{MR}}=0)}{\sqrt{S(f_{\text{MR}}=0)+B}}=3\,.$ (10)
-One can see that a fraction of magnetorotational collapses smaller than $7\%$ could be measured at $3\sigma$ after $20$ yr of data taking, assuming that the fraction of neutrino-driven BH-forming collapses is known and the neutral current atmospheric background is efficiently tagged.
+$\frac{S(f_{\text{MR}})-S(f_{\text{MR}}=0)}{\sqrt{S(f_{\text{MR}})+B}}=3\,.$ (10)
+One can see that a fraction of magnetorotational collapses smaller than $13\%$ ($19\%$) could be measured at $3\sigma$ after $20$ yr (10 yr) of data taking, assuming that the fraction of neutrino-driven BH-forming collapses is known and the neutral current atmospheric background is efficiently tagged.
 Figure 7: Hyper-Kamiokande sensitivity at $3\sigma$ to the fraction of magnetorotational core collapses ($f_{\text{MR}}$) as a function of the exposure.
@@
 The dark shaded band represents the impact of the uncertainties on the CCSN rate and the fraction of neutrino-driven BH-forming collapses.
-After $20$ yr exposure, Hyper-Kamiokande-Gadolinium will be sensitive to $f_{\rm{MR}}\gtrsim 7\%$.
+After $20$ yr exposure, Hyper-Kamiokande-Gadolinium will be sensitive to $f_{\rm{MR}}\gtrsim 13\%$
 The upcoming liquid Argon neutrino experiment DUNE [75], being sensitive to the $\nu_{e}$ component, could also add positively to this picture, although with smaller statistics (cf.
@@
 [83].
+Note that these sensitivity prospects rely on optimistic detector characteristics (e.g.
+perfect detection efficiency and a detector configuration foreseeing four tanks of $10$ kt each, whereas only the deployment of two tanks is envisioned for DUNE Phase I [84]).
 The difference between the mean energies of electron and non-electron flavor neutrinos is larger than the one for antineutrinos.
 This could potentially disentangle some of the existing uncertainties in the picture, especially for what concerns the flavor conversion physics and the fraction of neutrino-driven BH-forming collapses [29].
+However, even under optimistic assumptions, the contribution of DUNE to this matter is likely to be marginal.
 VI Complementary electromagnetic constraints
@@
 [40], to investigate the related neutrino properties and the associated contribution to the DSNB.
-Our models predict large neutrino luminosities $\gtrsim 10^{52}$ ergs-1 for several seconds and large mean energies for non-electron neutrinos and antineutrinos ($\langle\varepsilon_{\nu_{x}}\rangle\gtrsim 20$ MeV).
+Our models predict large neutrino luminosities $\gtrsim 10^{52}$ $\cdot$ s-1 for several seconds and large mean energies for non-electron neutrinos and antineutrinos ($\langle\varepsilon_{\nu_{x}}\rangle\gtrsim 20$ MeV).
 Such neutrino properties suggest that magnetorotational core collapses could provide a non-negligible contribution to the high-energy DSNB tail (above $5$ MeV).
+In this work, the evolution of the magnetorotational core collapses forming a protomagnetar is computed up to several seconds.
+A longer simulation would be desiderable to determine the evolution of the neutrino emission properties and assess the impact of the late-time emission on the diffuse emission and the potential numerical influence on the gradual rise of the mean energies at late times.
+The approach here adopted ensures that our results are conservative and do not overpredict the contribution of protomagnetars to the diffuse flux from MR core collapses.
 […]
-Given its larger size, the upcoming Hyper-Kamiokande-Gadolinium neutrino telescope could measure a fraction of magnetorotational collapses larger than $7\%$ at $3\sigma$ after $20$ yr of data-taking, if meanwhile we can infer the fraction of (neutrino-driven) BH-forming collapses and pin down the uncertanties on the CCSN rate through electromagnetic observations.
+Given its larger size, the upcoming Hyper-Kamiokande-Gadolinium neutrino telescope could measure a fraction of magnetorotational collapses larger than $13\%$ at $3\sigma$ after $20$ yr of data-taking, if meanwhile we can infer the fraction of (neutrino-driven) BH-forming collapses and pin down the uncertainties on the CCSN rate through electromagnetic observations.
 This measurement would further benefit from improved background reduction and detection efficiency.